How would i put x^2+4x+8=5x+18 in the standard ax^2+bx+c=0 form
Answer by TimothyLamb(4379)

---
given:
x^2 + 4x + 8 = 5x + 18
---
put in standard form:
---
x^2 + 4x + 8 - 5x - 18 = 0
---
answer:
x^2 - x - 10 = 0
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php